OpenStudy (anonymous):

in a sequence the sum of first n terms is given by Sn = nP+(1/2)n(n-1)Q, where P and Q are constants. Show that the sequence is AP. Find the first term, common difference and the 100th term. According to the answer given first term is P common difference is Q 100th term is P+99Q I am finding Tn= Sn+1-Sn = P + nQ but if we put T(1) then first term is coming P+Q T(100) = P+100 Q Is the answer wrong or I did the sum wrongly ??

OpenStudy (cwrw238):

plug in n = 1 and calculate the result this will give you the first term

OpenStudy (anonymous):

so then the first term would be P+Q !! but the answer given is P (the first term)

hartnn (hartnn):

after a lot of thought, i realized that Tn = Sn+1 - Sn is incorrect! it'd be Tn = Sn - S (n-1)

hartnn (hartnn):

Sum of n terms - sum of n-1 terms gives you n'th term., logical right ?? now you'll get your required result :)

OpenStudy (anonymous):

\[Sn-Sn-1 = nP-\frac{ 1 }{2 }n^2Q-\frac{ 1 }{2 }nQ-[(n-1)p+\frac{ 1 }{2 }(n-1)(n-2)Q]\]

hartnn (hartnn):

+1/2 n^2 Q

hartnn (hartnn):

you got that error ? it should be +1/2 n^2Q instead of -1/2n^2Q you wrote...

OpenStudy (anonymous):

thanks @hartnn for pointing out the mistake !! i got the answer !!

hartnn (hartnn):

welcome ^_^